Together, John and mark can finish painting a house in 10 hours. Alone, John takes 4 hours longer than Mark to finish painting the same house. How long does Mark take to finish painting the house by himself? Give your answer correct to 2 decimal place.
Answer by stanbon(75887)

Together, John and mark can finish painting a house in 10 hours.
Alone, John takes 4 hours longer than Mark to finish painting the same house. How long does Mark take to finish painting the house by himself?
----
Mark:: time = x hrs/job ; rate = 1/x job/hr
John:: time = x+4 hrs/job ; rate = 1/(x+4) job/hr
---
Together:: time = 10 hrs/job ; rate = 1/10 job/hr
-------------------
Equation:
rate + rate = together rate
1/x + 1/(x+4) = 1/10
------
10(x+4) + 10x = x(x+4)
------
20x + 40 = x^2 + 4x
------
x^2 - 16x - 40 = 0
-----
x = [16 +- sqrt(256-4*-40)]/2
-----
x = [16 +- sqrt(96)]/2
-----
x = [16 +- 4sqrt(6)]/2
----
x = [8 +- 2sqrt(6)]
-----
Comment: Strange answer for this type of problem.
Did you post it correctly?
Cheers,
Stan H.